Determine the y-intercept, the zeros, the number of turning points, and the behavior of the graph and sketch the graph of the following polynomial functions.


1. P(x) = x3 + 6x2 + 11x + 6


1
Expert's answer
2021-02-01T08:25:33-0500

y-intercept: "x=0"


"y(0)=(0)^3+6(0)^2+11(0)+6=6"

Point (0,6)

x-intercepts: "y=0"


"x^3+6x^2+11x+6=0"

"x^3+x^2+5x^2+11x+6=0"

"x^2(x+1)+(x+1)(5x+6)=0"

"(x+1)(x^2+5x+6)=0"

"(x+1)(x+2)(x+3)=0"

"x_1=-3, x_2=-2, x_1=-1"

There are 3 zeros.

Point (-3,0), point(-2, 0), and point(-1, 0).


"y'=3x^2+12x+11"

"y'=0:3x^2+12x+11=0"

"D=(12)^2-4(3)(11)=12"

"x_1=\dfrac{-12-\sqrt{12}}{2(3)}=-2-\dfrac{\sqrt{3}}{3}"

"x_2=\dfrac{-12+\sqrt{12}}{2(3)}=-2+\dfrac{\sqrt{3}}{3}"

There are 2 turning points.

"x<-2-\dfrac{\sqrt{3}}{3}, y'>0, y\ increases"

"-2-\dfrac{\sqrt{3}}{3}<x<-2+\dfrac{\sqrt{3}}{3}, y'<0, y\ decreases"

"x>-2+\dfrac{\sqrt{3}}{3}, y'>0, y\ increases"

The function has a local maximum at "x=-2-\dfrac{\sqrt{3}}{3}."

The function has a local minimum at "x=-2+\dfrac{\sqrt{3}}{3}."

Sketch the graph